Report of Activities on the RVIB N.B. Palmer
10 August 2001

Another long day at a deep water station. Work at station #47 took nearly 12 hours and we only recently have gotten underway for station #48, a station located on the continental shelf edge in 433 m of water. The heavy cloud layer persists with varying amounts of snow falling more or less continuously until recently when it turned into a light rain. Accumulations on the unheated decks amounted to a couple of inches. Currently on 11 August at 1435, the wind is out of the northwest (319) at 27 to 33 kts. The air temperature is a balmy 0.8°C (yes, above freezing), but the barometer is pretty low at 959.8 (a cold blast is likely to come our way next). Our position is -67° 19.518S; -74° 12.69W.

Yesterday (10 August), we completed work at three stations, #44, 45, and 46. The work was done under weather conditions that had not changed appreciably from the day before. A large low pressure system was still approaching and the barometer was slowly dropping. Relatively warm conditions with air temperatures around the freezing mark, strong winds in the 20 to 30 kt range, heavy dark clouds, and off and on snow prevailed. Included in the work were 3 CTDs, a 10-m2 MOCNESS trawl, a Tucker Trawl, one phytoplankton ring net tow. During the day, two sonabuoys were deployed. Unfortunately, although there were transits between stations during the day light period, the very poor visibility resulting from the falling and blowing snow prevented the bird and mammal surveyors from conducting their quantitative surveys. Only incidental observations were possible. BIOMAPER-II remained under repair and was not towyoed during this day.

John Klinck reports that on 10 August, the CTD group did three stations on the offshore end of the 300 line. Some concerns about leaking bottles are addressed. Casual ADCP observations look promising.

A number of leaking bottles has been detected from scatter in nutrient measurements which are taken from every bottle. Some of this leakage is from valves and some from bottles that do not seal. Some of the valves are difficult to seat or operate. A number of valves were fixed (by changing o-rings) before station #46. Two large o-rings were also replaced. We will see if this improves data quality.

Casual observation of the ADCP screen during CTD casts indicates that the surface layer often moves in a different direction from the rest of the visible water column (down to 300 m). The flow is strongly two layered with a very sharp transition at the base of the mixed layer. During a recent station at the shelf break, the water below the mixed layer was moving NE at about 15 cm/s, which is appropriate for our view of the ACC along this coast. Surface waters were moving SW at about 10 cm/s. Analysis of good ADCP records is likely to be a useful exercise.

Station #44 (399 m) is in the middle of the shelf with a 50 m deep mixed layer. The pycnocline extends over the next 100 m. A deep O2 minimum is seen at around 400 m, but there is no deep temperature maximum above the bottom. The warmest temperatures (1.38°C), on the bottom, indicate that this is well adjusted shelf water, long separated from the ACC.

Station #45 (384 m) is on the outer shelf and has many characteristics of station 44. Differences include a deeper mixed layer (65 m), a ragged pycnocline indicating active layering and a deep temperature maximum (1.48°C) at 270 m. The O2 minimum is around 250 m. The warmer deep temperatures shows influence from oceanic waters, which may be due to the nearness of this station to the shelf break.

Station 46 (1980 m) is off the shelf break on the 300 line and has a distinctly oceanic character. The mixed layer extends to about 80 m with a large warm intrusion just below causing considerable structure. The O2 minimum is about 250 m while a ragged temperature maximum (1.8°C) extends between 300 and 400 m. Below 500 m, the temperature decreases, and the salinity is nearly uniform.

Ana Sirovic reports that on 10 August, she deployed 2 difar sonobuoys. The first one was deployed as the MOC-10 was being towed on station #46, and she was able to monitor it for 1 h 19 min - until the signal was suddenly lost. It was clearly still not out of range, so the sonobuoy must have suffered damage from the ice. A second sonobuoy was deployed 11 nm from station #47. It was monitored for 1 hour. During this deployment, Jeff Otten was testing 2 preamplifiers on the Yagi antenna. Ana and Jeff decided that the preamps are not working properly when installed before the receiver in the dry lab. Also the GUV interference on the Yagi was significant, so the Sinclair omnidirectional antenna was used for this deployment. The Yagi problem will be worked on again when the weather improves and climbing up on the science mast sounds like a good idea. The signal from second buoy was lost at 6 nm away from the deployment site. No biological noise was heard on either deployment.

BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
Acoustical and optical surveying along the broad-scale trackline continued to be side-lined by the electronic problems with the SAIL communications loop. Although it appeared that the problems associated with the failure of the ESS data acquisition system had been identified and corrected, when the towed body was deployed for the transit to between stations #46 and 47, the system failed to work once in the water. BIOMAPER-II was again recovered and moved to the garage van, and another series of tests were conducted. This testing resulted in the replacement of a connector to the circuit board with the SAIL electronics. By the time the towed body was ready to enter the water, we had arrived at station #47 and only a test of the system in the water while on station was possible. This time the system worked and was ready for the towyoing along transect line # 7 back into Marguerite Bay.

No 1-m2 Mocness tows were taken on 10 August.

Cheers, Peter